Equivalent Units of Production and Related Costs

The charges to Work in Process—Baking Department for a period as well as information concerning production are as follows. The Baking Department uses the average cost method, and all direct materials are placed in process during production.

Work in Process—Baking Department
Bal., 1,700 units, 60% completed 3,774 To Finished Goods, 12,200 units ?
Direct materials, 12,400 units 40,300
Direct labor 19,667
Factory overhead 10,131
Bal., 1,900 units, 40% completed ?

If required, round cost per equivalent unit answer to two decimal places.

a. Determine the number of whole units to be accounted for and to be assigned costs.
units

b. Determine the number of equivalent units of production.
units

c. Determine the cost per equivalent unit.
$ per equivalent unit

d. Determine the cost of the units transferred to Finished Goods.
$

e. Determine the cost of units in ending Work in Process.
$
Answer #1

A) No of units to be accounted for = Beg units + WIP

= 1700+12400 = 14100 units

No of units Assigned for = Finished goods + closing stock

= 12200+1900 = 14100 units

B) No of Equivalent unit of production = Finished Goods + Closing Stock (completed)

= 12200+1900*.4

= 12960 units

C) Cost per Equivalent Units

= Opening cost + Direct material + Direct Labour + Factory Overhead / Total Equivalent units

= 3774 + 40300 + 19667 + 10131/12960

= $ 5.7/unit

D) Cost of units transferred to Finished goods

= 12200*5.7 = $69540

E) Cost of Units in Ending WIP

= 760*5.7 = $4332